OpenStudy (anonymous):

what is 2 + 2 + 1 - 1 =

OpenStudy (anonymous):

no its 4

OpenStudy (anonymous):

\[2 + 2 + 1 - 1 \rightarrow 2 + 2 = 4 \rightarrow 4 + 1 = 5 \rightarrow 5 - 1 = 4\]

OpenStudy (anonymous):

Calculator? You are on a computer...